OpenStudy (anonymous):

A family of 5 comes to the amusement park. There are two adults and three children (under the age of 12). They qualify for a special family rate of $174.45. The clerk tells them that a child's ticket always costs $11 less than an adult's ticket. Write an expression to represent the cost of the child's ticket in terms of the adult ticket. Remember the child's ticket is $11 less than the adult's ticket.

OpenStudy (anonymous):

let a = the cost of an adult ticket, then a-11 is the cost of a child's ticket (the way the question is phrased, this is the answer) this is an equation that represents the family situation 2a+3(a-11)=174.45 since there are 2 adults that pay $a and three children that pay $(a-11) each

OpenStudy (anonymous):

so the last equation represents the cost of a childs ticket

OpenStudy (anonymous):

a-11 is the cost of a child's ticket written in terms of the cost, a, of an adults ticket. if you are writing the answer, you should write: let a = the cost of an adult ticket, then a-11 is the cost of a child's ticket

OpenStudy (anonymous):

thanks this is part of a big project mind if i ask a few questions?

OpenStudy (anonymous):

soif for example an adults ticket cost $20 then a childs will be 11 dolars less, or 20-11=$9

OpenStudy (anonymous):

shoot

OpenStudy (anonymous):

ok kool i thing i get it it seems pretty easy so far but im not good at math

OpenStudy (anonymous):

2a+3(a-11)=174.45 i have to solve this wud it b 2a+3a-33=174.45 5a-33=174.45 +33 +33 5a=207.45 / 5 x=41.45?

OpenStudy (anonymous):

a=41.49 (not 45)

OpenStudy (anonymous):

yes my mistake but the rest is correct?? wow i feel kinda smart haha ok ummm i think i have like 2 more questions

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

ok and tht wud be the value of an adult ticket?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

so all i have to do wud be subtract 11 and tht wud be tha childs ticket

OpenStudy (anonymous):

yepper

OpenStudy (anonymous):

yaaay haha thnx

OpenStudy (anonymous):

you are kinda smart in math!

OpenStudy (anonymous):

hahaha yea i guess...not really :/

OpenStudy (anonymous):

umm ihave a part two i just asked the question wud u mind working with me to get it?

OpenStudy (anonymous):

ok